John likes coffee and cookies. The price for one cup of coffee is $1; the price for one cookie is $1. His marginal utility for coffee is 10 – x, where x is the amount of coffee. His marginal utility for cookies is 21 – 2y, where y is the number of cookies. How many cups of coffee and units of cookies will John buy if he has only $10?
1
Expert's answer
2020-08-21T11:31:40-0400
"x+y=10"

"\frac {10-x}{1}=\frac {21-2y}{1}"

"x=10-y"

"y=7"


"x=3"